OpenStudy (anonymous):

how do i Find the values of y = w(x) = square root of x for x= 0, 0.04, 0.09, 0.16, 0.25, 0.36, 0.49, 0.64, 0.81, 1. Then plot the corresponding points on a graph.

OpenStudy (freckles):

\[\text{ Example } .04=\frac{4}{100}=\frac{2^2}{10^2}=(\frac{2}{10})^2=(\frac{1}{5})^2\] Do you know the square root of (1/5)^2?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

1/25

OpenStudy (freckles):

square root and square cancel each other out it is 1/5

OpenStudy (anonymous):

ok

OpenStudy (freckles):

can you try the others now?

OpenStudy (anonymous):

ok i think i get it now